a rectangular plot has been fenced with 300 meters of long wire . find the length and breadth of the plot,if the length is twice its width. Also find the perimeter
Answer by Alan3354(69443)

a rectangular plot has been fenced with 300 meters of long wire. find the length and breadth of the plot,if the length is twice its width. Also find the perimeter
--------------
Find the perimeter? You just said it's 300 meters.
-------------
Anyway,
L = 2*W EQN 1
2L+2W = 300 EQN 2
-----------
Sub L in EQN 2
2*(2W) + 2W = 300
6W = 300
W = 50
----------
L = 100